Main Logbooks page LST commissioning Shift checklists Safe state checklists Maintenance activities LST2-4 Commissioning Old logbooks
GENERAL CAMERA DRIVE ENERGY AMC
  LST1 Drive ELOG  ELOG logo
Message ID: 265     Entry time: Tue Jul 9 00:46:52 2019
Author: Mitsunari, Yukiho, Noah, Moritz 
Type: Telescope Operations 
Category: Issues 
Subject: Park-in problems 
Status:  

At parking in, the telescope got stuck at 94.254°. We checked in the container that fuse Q21 went off. We did not see any errors both in the Drive cointainer and the GUI. We texted the experts. After 15 minutes, we pressed again "ParkIn" (without having touched the fuse). After nothing happend in the GUI, we called the experts. They authorized us to switch back the fuse. After that, still heavy intervention by the experts was needed, who finally made the telescope successfully park (fuse was on during all the trials). Thanks to the experts solving the issue at late night!

ELOG V3.1.5-fc6679b